### v2

# Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models

Abstract
Assessing the effectiveness of large language models (LLMs) in addressing diverse tasks is essential for comprehending their strengths and weaknesses. Conventional evaluation techniques typically apply a single prompting strategy uniformly across datasets, not considering the varying degrees of task complexity. We introduce the Hierarchical Prompting Taxonomy (HPT), a taxonomy that employs a Hierarchical Prompt Framework (HPF) composed of five unique prompting strategies, arranged from the simplest to the most complex, to assess LLMs more precisely and to offer a clearer perspective. This taxonomy assigns a score, called the Hierarchical Prompting Score (HP-Score), to datasets as well as LLMs based on the rules of the taxonomy, providing a nuanced understanding of their ability to solve diverse tasks and offering a universal measure of task complexity. Additionally, we introduce the Adaptive Hierarchical Prompt framework, which automates the selection of appropriate prompting strategies for each task. This study compares manual and adaptive hierarchical prompt frameworks using four instruction-tuned LLMs, namely Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B, across four datasets: BoolQ, CommonSenseQA (CSQA), IWSLT-2017 en-fr (IWSLT), and SamSum. Experiments demonstrate the effectiveness of HPT, providing a reliable way to compare different tasks and LLM capabilities. This paper leads to the development of a universal evaluation metric that can be used to evaluate both the complexity of the datasets and the capabilities of LLMs. The implementation of both manual HPF and adaptive HPF is publicly available [^1] GitHub repository consisting of the code implementation of this work: https://github.com/devichand579/HPT.
Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models
Devichand Budagam1, Sankalp KJ2, Ashutosh Kumar3, Vinija Jain4, 5thanks: Work does not relate to position at Amazon, Aman Chadha4, 5thanks: Work does not relate to position at Amazon 1Indian Institute of Technology Kharagpur, India, 2AI Institute, University of South Carolina, USA 3Rochester Institute of Technology, USA, 4Amazon GenAI, USA, 5Stanford University, USA
1 Introduction
Large language models (LLMs) have revolutionized natural language processing (NLP), enabling significant advancements across a wide range of applications. However, evaluating the capabilities of LLMs across various datasets remains a challenging task. Traditional prompting strategies, which apply the same prompt to all dataset samples regardless of task complexity, often result in sub-optimal evaluations. This limitation highlights the need for more sophisticated evaluation methods that can adapt to varying levels of task complexity.
Figure 1: Hierarchical Prompt Framework (HPF) consists of five distinct prompting strategies, each tailored to different levels of task complexity, ensuring the appropriate prompt is selected for the given task.
Prompting plays a vital role in unlocking the full potential of LLMs. By designing specific input prompts, the LLM’s responses can be guided, significantly influencing the quality and relevance of the output. Effective prompting strategies have enhanced LLM performance on tasks ranging from simple question-answering to complex reasoning and problem-solving. Recent research has explored various approaches to prompting and reasoning evaluation in LLMs. Chain-of-Thought (CoT) prompting Wei et al. (2022b) elicits step-by-step reasoning, improving performance on complex tasks. Specializing smaller models Fu et al. (2023) and using large models as reasoning teachers Ho et al. (2022) have demonstrated the potential for enhancing reasoning capabilities. Progressive-hint prompting Zheng et al. (2023) has been proposed to conceptualize answer generation and guide LLMs toward correct responses. Metacognitive prompting Wang and Zhao (2024) incorporates self-aware evaluations to enhance understanding abilities.
We introduce the Hierarchical Prompting Taxonomy (HPT), a set of rules that employs a Hierarchical Prompt Framework (HPF) shown in Figure 1, a structured evaluation method addressing the need for more sophisticated assessment of LLM task-solving capabilities. This framework employs five distinct prompting strategies, each tailored to different levels of task complexity. By assigning a score called the HP-Score based on the LLM’s performance across these levels, this work provides a comprehensive assessment of its ability to solve diverse tasks.
Furthermore, the taxonomy incorporates an Adaptive Hierarchical Prompt framework, enabling an LLM, called prompt-selector to dynamically select the appropriate prompting strategy based on a given task complexity. This adaptive approach automates the evaluation process, also provides a better understanding of the judgment skills of LLMs in determining task complexity. The main contributions of this paper are as follows:
  1. 1.
    We propose the Hierarchical Prompting Taxonomy (HPT), a set of rules designed to establish a universal measure of task complexity for both datasets and LLMs.
  2. 2.
    Based on the rules of the HPT, this work introduces a set of five unique prompting strategies, arranged from simplest to most complex, to provide a more precise assessment of LLMs and offer clearer insights.
  3. 3.
    We introduce the Adaptive HPF that allows us to examine the judgement skills of LLMs by automating the evaluation process and dynamically selecting the right prompting approach based on task complexity.
The rest of the paper is organized as follows: Section 2 provides an overview of related work on prompting and evaluation in LLMs. Section 3 presents HPT and the frameworks associated with it in detail. Section 4 describes the experimental setup and results. Section 5 concludes the paper, Section 6 discusses the limitations and finally, Section 7 discusses the ethical impact of this work.
2 Related Work
The advent of LLMs has revolutionized NLP by demonstrating significant improvements in few-shot and zero-shot learning capabilities. Brown et al. (2020) introduced GPT-3, a 175 billion parameter autoregressive model, showcasing its ability to perform a wide range of tasks such as question-answering, reading comprehension, translation, and natural language inference without fine-tuning. This study highlighted the potential of very large models for in-context learning, while also identifying limitations in commonsense reasoning and specific comprehension tasks. Similarly, Liu et al. (2021) surveyed prompt-based learning, emphasizing the role of prompt engineering in leveraging pre-trained models for few-shot and zero-shot adaptation to new tasks with minimal labeled data.
Emergent abilities in LLMs, which appear suddenly at certain scale thresholds, have also been a topic of interest. Wei et al. (2022a) examined these abilities in few-shot prompting, discussing the underlying factors and implications for future scaling. Complementing this, Kojima et al. (2022) demonstrated that LLMs could exhibit multi-step reasoning capabilities in a zero-shot setting by simply modifying the prompt structure, thus highlighting their potential as general reasoning engines. Further exploring reasoning capabilities, Qiao et al. (2023) categorized various prompting methods and evaluated their effectiveness across different model scales and reasoning tasks, identifying key open questions for achieving robust and generalizable reasoning.
Several studies have focused on improving the interaction and decision-making abilities of LLMs through innovative prompting techniques. Lo et al. (2023) proposed ASH (Actor-Summarizer-Hierarchical) prompting, which enhances LLM performance on complex web navigation tasks by decomposing the process into summarization and action prediction. This hierarchical approach significantly outperformed previous methods on tasks with lengthy trajectories. Kong et al. (2024) introduced role-play prompting to improve zero-shot reasoning by constructing role-immersion interactions, which implicitly trigger chain-of-thought processes and enhance performance across diverse reasoning benchmarks.
The robustness and safety of LLM outputs have also been addressed in recent research. Wang et al. (2023) developed INDUST, a benchmark to evaluate LLM resistance to inductive instructions containing false premises, showing that self-critique prompting can help models handle such challenging instructions more effectively. Additionally, Chen et al. (2024) provided a comprehensive survey on prompt engineering techniques aimed at reducing hallucinations and enhancing the reliability of LLM outputs through advanced methods like chain-of-thought and tree-of-thought prompting.
To facilitate the evaluation and understanding of LLM capabilities, Zhu et al. (2024) introduced PromptBench, a unified library encompassing a variety of LLMs, datasets, evaluation protocols, and adversarial prompt attacks. This modular and extensible tool aims to support collaborative research and advance the comprehension of LLM strengths and weaknesses. Zhou et al. (2023b) proposed DYNAICL, a framework for efficient prompting that dynamically allocates in-context examples based on a meta-controller’s predictions, achieving better performance-efficiency trade-offs compared to uniform example allocation.
Chang et al. (2022) proposed SpeechPrompt, a framework that reformulates tasks as conditional generation problems for pre-trained generative spoken language models. This method demonstrated competitive performance in speech classification with fewer trainable parameters and potential in more challenging tasks like automatic speech recognition.
Lastly, Yao et al. (2023) introduced the Tree-of-Thoughts framework, enabling LLMs to deliberate over coherent text units and perform heuristic searches for complex reasoning tasks. This approach generalizes over chain-of-thought prompting and has shown significant performance improvements in tasks requiring planning and search, such as creative writing and problem-solving games.
These works collectively highlight the advancements in leveraging LLMs through innovative prompting techniques, addressing their emergent abilities, reasoning capabilities, interaction strategies, robustness, and evaluation methodologies. Despite significant advancements, the current research on LLMs reveals several limitations, particularly in terms of prompt design, handling complex reasoning tasks, and evaluating model performance across diverse scenarios. The emergent abilities of LLMs, while promising, often lack predictability and control, and the robustness of these LLMs in the face of misleading prompts remains a concern.
To address these challenges, we propose the HPT and HPF for evaluating different LLMs. This approach aims to systematically categorize and assess the effectiveness of various prompting techniques based on task complexity, enabling a more nuanced understanding of LLM behaviors and performance. By introducing a structured framework for evaluation, this work seeks to enhance the reliability and applicability of LLMs, facilitating their deployment in increasingly complex and dynamic environments.
3 Methodology
Figure 2: Hierarchical Prompting Taxonomy (HPT): A taxonomy designed to assess task complexity based on the criteria: Basic Recall and Reproduction, Understanding and Interpretation, Analysis and Reasoning, and Application of Knowledge and Reasoning.
3.1 Datasets and Models
Experiments were performed on BoolQ Clark et al. (2019), CommonSenseQA (CSQA) Talmor et al. (2019), IWSLT-2017 en-fr (IWSLT) Cettolo et al. (2017), and SamSum Gliwa et al. (2019). BoolQ consists of approximately 16,000 True/False questions with corresponding passages from Wikipedia, designed for binary classification in reading comprehension. CSQA contains around 12,000 multiple-choice questions sourced from ConceptNet, aimed at evaluating model’s commonsense reasoning abilities. IWSLT is a parallel corpus with tens of thousands of English-French sentence pairs from TED Talks, used for machine translation tasks. SamSum features around 16,000 human-generated chat logs with summaries, facilitating dialogue summarization model development. On the aforementioned datasets, the performance of the instruction-tuned LLMs Llama 3 8B AI@Meta (2024), Phi 3 3.8B Abdin et al. (2024), Mistral 7B Jiang et al. (2023), and Gemma 7B Team et al. (2024) were evaluated on their validation splits, with the exception of SamSum, which utilized the test split and nearly 2300 hours of computation was used for all experiments on the V100 GPU.
3.2 Hierarchical Prompting Taxonomy (HPT)
The proposed HPT shown in Figure 2 is founded on principles that are crucial for humans to address the majority of reasoning and interpretation tasks, taking into account the following criteria:
  • Basic Recall and Reproduction: Remember and reproduce factual information without interpretation or analysis.
  • Understanding and Interpretation: Comprehend and explain the meaning of information, summarizing or clarifying content.
  • Analysis and Reasoning: Break down complex information, understand relationships, and solve problems using logical reasoning.
  • Application of Knowledge and Execution: Apply knowledge in practical situations, execute multi-step processes, and solve complex tasks.
3.3 Hierarchical Prompt Framework (HPF)
The HPT employs an HPF that comprises five different prompting strategies. The complexity levels of these prompting strategies within the framework are determined by the taxonomy criteria, which judges actions such as the requirement for external knowledge, the capacity to solve subproblems, and the capacity to provide step-by-step reasoning that takes the task’s complexity into account. The five different prompting strategies that make up the HPF are :
  • Role Prompting Kong et al. (2024): Prompts that define a role for the LLM in solving the task.
  • Zero-Shot Chain-of-Thought Prompting (Zero-CoT) Kojima et al. (2022): Prompts that use the phrase "Let’s think step by step" without providing prior examples to encourage critical thinking and problem-solving.
  • Three-Shot Chain-of-Thought Prompting (3-CoT) Wei et al. (2022b): Prompts that offer three examples to guide the LLM’s reasoning process.
  • Least-to-Most Prompting Zhou et al. (2023a): Prompts that sequentially breakdown the task into sub problems to derive essential insights from the task in order to solve it.
  • Generated Knowledge Prompting (GKP) Liu et al. (2022): Prompts that require the integration of external knowledge to accomplish the task. Llama 3 8B was used to generate the necessary knowledge. Unlike the original study, the knowledge prediction with the highest confidence score was selected among three possible predictions.
The Manual HPF systematically iterates through the predefined levels of prompting strategies to solve a given task using the language models. This process involves starting with the simplest prompts and gradually progressing to more complex ones based on the prompt hierarchy levels until the task is successfully addressed. By manually traversing each level, it ensures that every potential strategy is explored and allowing the LLM to identify the most effective approach for the task.
A prompt-selector determines which prompting strategy is most suited for a given task in the adaptive HPF. This allows us to examine how LLMs may be utilised to assess their abilities to do different tasks while also automating the evaluation process.
3.4 Hierarchical Prompting Score (HP-Score)
The HP-Score is an evaluation metric for assessing the complexity of a task in comparison to various categories of problem-solving agents, such as humans and LLMs, which is heavily influenced by HPT rules. A task’s lower HP-Score indicates that the corresponding agent is more capable of solving the task.
3.4.1 HP-Score for Datasets
Datasets are evaluated by scoring each sample of a representative set of 5% of the original evaluation dataset on a scale from 1 (lowest) to the number of prompting levels (highest) existing in HPF for each of the four HPT rules. HP-Score for datasets is computed as:
$\text{hp}_{\text{j}}=\frac{1}{4}\sum_{i=1}^{4}a_{i}$ (1)
[…]
where $a_{i}$ represents the individual score for each rule of HPT, $\text{hp}_{\text{j}}$ represents the HP-Score of sample $j$ and $n$ denotes the total number of samples in the representative set. The HP-Scores assigned by experts in the domains related to the datasets used in the experiments are displayed in Table 1. CSQA is the most complex task relative to humans, with the highest HP-Score among the four datasets.
Dataset Evaluation Set Size Representative Set Size HP-Score
BoolQ 3270 162 1.71
CSQA 1221 60 2.52
IWSLT 890 45 1.92
SamSum 819 40 2.23
Table 1: HP-Scores of datasets by human experts.
3.4.2 HP-Score for Manual HPF
HPF categorizes prompting strategies based on their complexity into five levels: Role Prompting (level 1), Zero-CoT (level 2), 3-CoT (level 3), Least-to-Most Prompting (level 4), and GKP (level 5). Each level has a score of the same value as the level, i.e. for level x, score = x. A LLM’s superior task-solving ability is indicated by a lower HP-Score. If the LLM fails to address the task at all levels of prompts, it is assigned a penalty corresponding to its dataset HP-Score. If the LLM was able to address the task at level x:
$\text{hp}_{\text{j}}=x$ (3)
If LLM was unable to address the task after all m levels:
$\text{hp}_{\text{j}}=m+\text{HP-Score}_{\text{Dataset}}$ (4)
HP-Score for Manual HPF is:
[…]
where $\text{hp}_{\text{j}}$ represents the HP-Score of sample $j$, x represents the level of HPF at which task is addressed, $m$ represents the total number of levels in HPF, and $n$ denotes the total number of samples in the evaluation set.
3.4.3 HP-Score for Adaptive HPF
The prompt-selector can dynamically select the most suitable prompting strategy for a given task’s complexity from the manual HPF’s collection of prompting strategies. In order to determine the most effective prompting strategy to address the task, the prompt-selector was given a maximum number of iterations equivalent to the number of levels in the manual HPF. The score for ith iteration is i + x, where x is the level of the prompting strategy selected by the prompt-selector. If the LLM fails to address the task after all iterations, it is assigned a penalty corresponding to it’s dataset HP-Score. If LLM was able to address the task at ith iteration:
$\text{hp}_{\text{j}}=x+i$ (6)
If LLM was unable to address the task after all m iterations:
$\text{hp}_{\text{j}}=m+\text{HP-Score}_{\text{Dataset}}$ (7)
HP-Score for Adaptive HPF is:
[…]
where $\text{hp}_{\text{j}}$ represents the HP-Score of sample $j$, x represents the level of HPF selected by prompt-selector at $i$th iteration at which task is addressed, $m$ represents the total number of levels in HPF, and $n$ denotes the total number of samples in the evaluation set.
4 Results
4.1 Results for Manual HPF
Table 2 presents the HP-Scores for the four LLMs: Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B, across previously mentioned datasets with varying threshold values (0.15 and 0.20) for BLEU Papineni et al. (2002) and ROUGE Lin (2004). The BLEU and ROUGE thresholds at every prompting level serve as the task-solving criteria for each sample of IWSLT and SamSum respectively. The complexity at which the LLMs are able to handle the dataset effectively is reflected in the HP-Scores.
Figure 3: HP-Scores of all datasets and threshold values for LLMs in Manual HPF.
Model BoolQ CSQA IWSLT (0.15) IWSLT (0.20) SamSum (0.15) SamSum (0.20)
Llama 3 8B 1.3692 2.4281 3.3955 3.9246 1.3018 1.7166
Mistral 7B 1.4063 2.4929 3.5183 4.1455 1.2590 1.6770
Phi 3 3.8B 1.3965 2.1714 3.8479 4.4401 2.0925 3.4010
Gemma 7B 1.4558 2.5053 5.3929 5.8369 3.3137 5.0316
Table 2: HP-Scores of datasets (with thresholds) across all LLMs for Manual HPF.
Dataset Metric Threshold Llama 3 8B Phi 3 3.8B Mistral 7B Gemma 7B
BoolQ Accuracy - 0.99296 0.98685 0.98073 0.99419
CSQA Accuracy - 0.84766 0.88452 0.82063 0.83783
IWSLT BLEU 0.15 0.23539 0.19327 0.21629 0.08049
0.2 0.23091 0.19098 0.21569 0.06807
SamSum ROUGE-1 0.15 0.29395 0.21389 0.28529 0.16067
0.2 0.29996 0.22109 0.29365 0.14063
Table 3: Evaluation scores for datasets across all LLMs for Manual HPF.
Figure 3 shows that Llama 3 8B achieves lower HP-Scores than other LLMs across all datasets, indicating superior performance and the ability to solve tasks with simpler prompting strategies. All four LLMs were better at solving BoolQ than humans, with lower HP-Scores than BoolQ HP-Score by human experts. IWSLT was the task where all four LLMs were incapable of solving even at threshold values of 0.15 and 0.20, indicating that LLMs are incapable in solving domain-specific tasks compared to human experts in the respective domains.
Table 3 shows the evaluation scores for each LLM across all datasets using the Manual HPF. Llama 3 8B and Gemma 7B outperform Phi 3 3.8B and Mistral 7B on BoolQ, with all four LLMs achieving near-perfect accuracy. Phi 3 3.8B outperformed all other LLMs in solving CSQA. Despite having the fewest parameters, phi 3 3.8B achieved the lowest HP-Score among all four LLMs while also achieving the highest accuracy. Llama 3 8B has the highest BLEU scores and ROUGE scores across both threshold values for IWSLT and SamSum datasets respectively. While Gemma 7B has the lowest performance and the highest HP-Scores for the IWSLT and SamSum datasets. Llama 3 8B and Mistral 7B are the most capable and genearlized LLMs, achieving superior performance across all datasets. Phi 3 3.8B, even with the fewest parameters, was able to compete with Llama 3 8B and Mistral 7B. Gemma 7B even though was able to solve BoolQ and CSQA with superior performance showing it’s reasoning abilities but was unable to achieve good performance across domain-specific datasets like IWSLT and SamSum.
Model BoolQ CSQA IWSLT (0.15) IWSLT (0.20) SamSum (0.15) SamSum (0.20)
Llama3 5.2173 5.9136 6.2006 6.2841 5.0316 5.5756
Mistra 7B 5.0483 5.9073 6.2478 6.4604 4.7423 5.1336
Phi 3 3.8B 5.1386 5.6793 6.3955 6.4936 5.0961 5.7778
Gemma 7B 5.1514 5.5771 6.5947 6.6605 5.7229 6.4347
Table 4: HP-Scores of datasets (with thresholds) across all LLMs for Adaptive HPF.
Dataset Metric Threshold Llama3 8B Phi 3 3.8B Mistral 7B Gemma 7B
BoolQ Accuracy - 0.88577 0.91115 0.91752 0.91166
CSQA Accuracy - 0.59451 0.68019 0.60111 0.68549
IWSLT BLEU 0.15 0.21140 0.15557 0.20000 0.08447
0.2 0.21146 0.15354 0.20568 0.07730
SamSum ROUGE-1 0.15 0.24407 0.20586 0.26910 0.16023
0.2 0.24981 0.21580 0.28335 0.16179
Table 5: Evaluation scores for datasets across all LLMs for Adaptive HPF.
4.2 Results for Adaptive HPF
The Adaptive HPF automates the selection of appropriate prompting strategy for each task using a prompt-selector utilizing Llama 3 8B for the experiments. Table 4 shows the HP-Scores and Table 5 presents the evaluation scores for the LLMs using the adaptive HPF.
Figure 4: HP-Scores of all datasets and threshold values for LLMs in Adaptive HPF.
The findings presented in Figure 4 show that, on all datasets, the adaptive HPF produces HP-Scores for the LLMs that are higher than those obtained with the manual HPF, with decreased evaluation scores. Higher HP-Scores demonstrate that adaptive HPF is obviously unable to complete a variety of tasks, with the primary cause being hallucinations when choosing the appropriate level of prompting by the prompt-selector. Hallucination of prompt-selector leads to repeatedly choosing irrelevant prompting levels or choosing levels out of the range of manual HPF for several iterations of adaptive HPF, resulting in high HP-Scores. This demonstrates that LLMs are incapable of judging their evaluations of various tasks.
Figure 5: Accuracy scores of BoolQ and CSQA datasets across all LLMs for manual HPF and Adaptive HPF.
Figure 6: BLEU scores of IWSLT dataset across all LLMs for manual HPF and Adaptive HPF..
Figure 7: ROUGE-1 scores of SamSum dataset across all LLMs for manual HPF and Adaptive HPF..
4.3 Comparison between Manual HPF and Adaptive HPF
Figure 5 illustrates that the manual HPF outperforms the adaptive HPF on the BoolQ dataset across all LLMs. The decrease in accuracy with the adaptive HPF demonstrates its inability to select the most appropriate prompting level within its iteration limit. Figure 5 shows a significant decline in performance for LLMs in solving CSQA, suggesting that manual HPF is more dependable than adaptive HPF for solving CSQA. Despite solving the IWSLT with slightly worse performance as shown in Figure 6 when using the adaptive HPF, LLMs are unable to identify the optimal prompting level indicated by their high HP-Scores. Figure 7 shows that both frameworks perform similarly in solving SamSum, with the adaptive HPF showing higher HP-Scores across all LLMs. Table 1 shows that the HP-Scores of both manual HPF and adaptive HPF indicate that manual HPF more closely matches human estimation of task complexity, while the adaptive HPF faces difficulty in solving these tasks when compared to both manual HPF and humans. Both frameworks have higher HP-Scores when solving IWSLT than domain experts’ HP-Scores, indicating that pre-trained LLMs with more domain-specific knowledge are needed to demonstrate better performance across domain-specific tasks with less complex prompting strategies, reducing the need for these LLMs to be fine-tuned. Another important finding from the IWSLT and SamSum evaluation scores is that all four LLMs were unable to perform above a particular evaluation score at higher threshold values, as evidenced by the saturation of their performance across all threshold values.
Overall, our experiments demonstrate that different tasks don’t always require sophisticated prompting strategies from LLMs; it largely depends on the LLM’s aptitude for the task. The HP-Score provides a universal measure of task complexity relative to different agents, offering insights into the capabilities of both LLMs and humans. When comparing Manual HPF and Adaptive HPF, the Manual HPF is more reliable and consistently performs well on datasets with lower HP-Scores. In contrast, the Adaptive HPF is computationally inefficient due to hallucinations when selecting the prompting level in least number of iterations. For other tasks and LLMs, Adaptive HPF requires finetuning of the prompt selector or other techniques to minimize hallucinations and ensure consistent outcomes.
5 Conclusion
The HPT, combined with the HPFs, offers a robust and efficient method for evaluating LLMs based entirely on task complexity. These findings highlight the effectiveness of the HPF in evaluating the task with the most effective prompting strategies based on task complexity, leading to enhanced LLM performance across various datasets and tasks. This approach not only provides a detailed understanding of LLMs problem-solving capabilities but also introduces Adaptive HPF by dynamically selecting the most appropriate prompting strategy, paving the way for future advancements in LLM evaluation techniques.
6 Limitations
Restricted Model Evaluation: The study focused on four specific instruction-tuned LLMs: Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B. While these LLMs provide valuable insights, the inclusion of a more diverse set of LLMs, both in terms of architecture and scale, would help assess the applicability of the HPT across a broader spectrum of LLMs.
Restricted Dataset Evaluation: The evaluation of both HPF and adaptive HPF is conducted on four distinct datasets encompassing tasks such as question-answering, reading comprehension, translation, and summarization. Incorporating a wider variety of datasets that cover additional tasks would aid in evaluating the HPT’s applicability across a more extensive range of tasks.
Manual framework optimization: The effectiveness of the HPF heavily relies on the quality of the prompts used at each level of the taxonomy. Crafting high-quality prompts that accurately reflect the subtleties of each level demands considerable expertise and repeated refinement. This study only investigated a limited set of prompting strategies within the HPF, indicating a need for further research into creating diverse structural frameworks and incorporating additional prompting strategies.
Adaptive framework optimization: The Adaptive HPF automates the selection of suitable prompting strategies according to task complexity. However, the current system depends on a prompt-selector that utilizes an LLM. There were numerous hallucinations when choosing the appropriate prompting strategy for the task. Further investigation is required to minimize hallucinations in the prompt-selector and to explore alternative methods, such as ensemble techniques or meta-learning approaches, to enhance the efficiency and effectiveness of the adaptive framework.
7 Ethical Consideration
The HP-Scores assigned by experts in the domains related to the datasets—BoolQ, CSQA, IWSLT and SamSum—may introduce bias into the comparative analysis. This potential bias stems from the subjective nature of expert scoring, which can be influenced by individual experience and perspective. Despite this, the datasets utilized in this study are publicly available and widely recognized in the research community, thereby minimizing the risk of unanticipated ethical issues. Nevertheless, it is crucial to acknowledge the possibility of scoring bias to ensure transparency and integrity in the analysis.
Appendix A Large Language Models Used for Evaluation
The HPF supports multiple LLMs and includes mechanisms for optimizing performance through advanced quantization techniques. The experiments were conducted on the following instruction-tuned LLMs and the model description and licenses are discussed in Table 6.
Model License Type Usage Restrictions
Llama 3 8B Research License Non-commercial, requires attribution, ethical guidelines
Phi 3 3.8B Open-source Broad use allowed, must include original license and notices
Mistral 7B Open-source Broad use allowed, must include original license and notices
Gemma 7B Open-source/Research Depending on license, may have non-commercial restrictions or broad use allowed
Table 6: License information for LLMs used in the experiments.
The selection of Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B offers a comprehensive evaluation set that spans different model sizes, architectures, and use cases, ensuring the HPF framework is robust and generalizable. Llama 3 8B excels in language understanding and generative tasks, achieving top-tier performance across benchmarks such as MMLU, HumanEval, GSM-8K, BoolQ, and CommonSenseQA. This model’s proficiency in diverse tasks, including complex mathematical reasoning and commonsense reasoning, underscores its high-quality text generation capabilities. Phi 3 3.8B, optimized for environments with limited computational resources, performs competitively on benchmarks like GSM-8K and BoolQ, making it ideal for real-time applications due to its efficiency and quick response times. Mistral 7B balances parameter count and performance, excelling in MMLU, CommonSenseQA, and HumanEval, showcasing versatility in language understanding and code-related tasks. Gemma 7B, specialized for particular domains, excels in niche areas like medical text understanding and technical documentation. While not always the top performer in general benchmarks, Gemma 7B’s specialization complements the general-purpose nature of the other LLMs, ensuring the HPF framework is applicable to a wide range of scenarios and tasks.
By including these LLMs, the effectiveness of the HPF is highlighted, demonstrating its ability to accommodate a wide range of scenarios and tasks, from general-purpose language understanding and generation to specialized domain-specific applications. This comprehensive evaluation ensures that HPF is a robust and generalizable framework in the landscape of LLM evaluation and prompting frameworks.
The quantization configuration for each of the instruction-tuned LLM involves reducing the precision of the LLM’s weights and computations, allowing for a more compact representation and faster processing. The LLMs were loaded in a 4-bit precision format which significantly reduces the memory footprint without compromising accuracy. By using half-precision floating point computations (float16), the LLMs maintain high computational efficiency. Additionally, the quantization type employed is NF4, which provides a balanced approach between model size and performance. The use of double quantization further refines this process, ensuring that the LLMs operate efficiently even on resource-constrained hardware.
The generation configuration is carefully designed to optimize the quality of text output during evaluation. This involves setting parameters that influence the behavior of the LLMs during the text generation process. For instance, the maximum number of tokens that can be generated in a single run is capped (1024 tokens) to ensure concise and relevant outputs. The temperature parameter was set to 0.6 to control the randomness of the predictions; lower temperature results in more deterministic and focused outputs. Top-p (p=0.9) sampling allows the LLMs to explore various possible continuations for a given input. Finally, the repetition penalty is applied to discourage the LLMs from repeating the same phrases or sentences, thus ensuring more coherent and varied text generation.
Appendix B Prompt Templates
B.1 Level 1: Role Prompting
Role prompting is the most basic level of interaction with a LLM, where the LLM is given a specific role or task to perform without additional context or examples. This technique relies solely on the initial instruction to guide the LLM’s response. For instance, if the LLM is asked to "act as a translator", it will attempt to translate text based on its training data. This method is straightforward but may lack depth and context, leading to less accurate or nuanced outputs. Table 7 presents the prompt templates utilized for Role prompting across all datasets in our experiments.
Dataset Prompt
BoolQ Based on the passage: ‘‘passage", answer True/False to the question: ‘‘question" as an Omniscient person.
CSQA Choose the answer: ‘‘question",A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5" as a critical thinker.
IWSLT Translate ‘‘english text" to french as a Translator.
SamSum Summarize the Dialogue: ‘‘dialogue" as a Summarizer.
Table 7: Prompt templates of different datasets for Role Prompting.
B.2 Level 2: Zero-shot Chain-of-Thought Prompting
Zero-shot Chain-of-Thought prompting builds upon basic role prompting by asking the LLM to generate a reasoning process for a given task, even though it has not been explicitly trained on similar examples. In this approach, the LLM attempts to break down the problem and solve it step-by-step using its internal knowledge base. This technique leverages the LLM’s ability to infer and reason, improving the quality of the response by encouraging logical progression and coherence in the thought process. Table 8 presents the prompt templates utilized for Zero-CoT across all datasets in our experiments.
Dataset Prompt
BoolQ Based on the passage: ‘‘passage", answer True/False to the question: ‘‘question". Let’s think step by step.
CSQA Choose the answer: A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5". Let’s think step by step.
IWSLT Translate ‘‘english text" to french. Let’s think step by step.
SamSum Summarize the Dialogue: ‘‘dialogue". Let’s think step by step.
Table 8: Prompt templates of different datasets for Zero-shot Chain-of-Thought Prompting.
B.3 Level 3: Three-Shot Chain-of-Thought Prompting
Three-shot Chain-of-Thought prompting advances the zero-shot approach by providing the LLM with three examples of the task at hand, including the reasoning steps taken to arrive at the solution. These examples serve as a guide, helping the LLM understand the structure and logic required to solve similar problems. By seeing how previous tasks were approached and resolved, the LLM can better mimic the process, leading to more accurate and contextually appropriate responses. Table 9 presents the prompt templates utilized for 3-CoT across all datasets in our experiments.
Dataset Prompt
BoolQ Based on the passage: "passage1", answer True/False to the question: "question1". Answer: "answer1". Explanation: "explaination1". Based on the passage: "passage2", Answer True/False to the question: "question2". Answer: "answer2". Explanation: "explaination2". Based on the passage: "passage3", Answer True/False to the question: "question3". Answer: "answer3". Explanation: "explaination3". Based on the passage: "passage", Answer True/False to the question: "question".
CSQA Choose the answer: "question1",A. "option1-1",B. "option2-1",C. "option3-1",D. "option4-1",E. "option5-1", Answer: "answer1", Explanation: "explaination1". Choose the answer: "question2", A. "option1-2",B. "option2-2",C. "option3-2",D. "option4-2",E. "option5-2", Answer: "answer2", Explanation: "explainatio n2". Choose the answer: "question3", A. "option1-3",B. "option2-3",C. "option3-3",D. "option4-3",E. "option5-3", Answer: "answer3", Explanation: "explaination3". Choose the answer: "question", ‘‘question",A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5".
IWSLT Translate "english text1" to French. French: "french text1". Translate "english text2" to French. French: "french text2". Translate "english text3" to French. French: "french text3". Translate "english text" to French.
SamSum Summarize the Dialogue: "dialogue1". Summary: "summary1". Summarize the Dialogue: "dialogue2". Summary: "summary2". Summarize the Dialogue: "dialogue3". Summary: "summary3". Summarize the Dialogue: "dialogue".
Table 9: Prompt templates of different datasets for Three-Shot Chain-of-Thought Prompting
B.4 Level 4: Least-to-Most Prompting
Least-to-most prompting is a sophisticated technique that involves gradually increasing the complexity of prompts presented to the LLM. The process begins with simpler, more manageable tasks and progressively moves to more complex challenges. This method helps the LLM build confidence and competence in solving problems, as it can leverage the insights gained from simpler prompts to tackle more difficult ones. It enhances the LLM’s ability to generalize from straightforward examples to more intricate scenarios. Table 10 presents the prompt templates utilized for Least-to-Most Prompting across all datasets in our experiments.
Dataset Prompt
BoolQ prompt 1: Summarize the main points of this passage: "passage". prompt 2: Analyze this question to identify its key components: "question".
prompt 3: Find the part of the passage that relates to this question: "question", Passage: "passage".
prompt 4: Based on the passage, what is the answer to this question: "question", Relevant Information: "previous response".
CSQA prompt 1: Analyze this question: "question".
prompt 2: Elaborate about each option for the question: "question", options: A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5".
prompt 3: Based on the analysis: "previous response", discard wrong answers among the options: A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5".
prompt 4: Choose the correct answer from the options: A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5".
IWSLT prompt 1: What is the main idea or theme of this text? "english text".
prompt 2: Identify and list the key phrases or terms in this text: "english text".
prompt 3: Translate the following key phrases into French: "previous response".
prompt 4: Translate "english text" into French, incorporating the translations of the key phrases: "previous response".
SamSum prompt 1: List the main points or key ideas present in this dialogue: "dialogue".
prompt 2: Elaborate on the following key points, providing additional details or context: "previous response".
prompt 3: Using the listed key points and their elaborations, draft a concise summary of this text: "dialogue".
prompt 4: Refine this draft summary to make it more concise and coherent, ensuring it captures the essence of the text: "dialogue".
Table 10: Prompt templates of different datasets for Least-to-Most Prompting
B.5 Level 5: Generated Knowledge Prompting
Generated Knowledge prompting represents one of the most complex prompting techniques present in the HPF, where the LLM not only addresses the given task but also incorporates additional relevant information to enhance its response. This method involves prompting another LLM to produce auxiliary knowledge that can be used to better understand and solve the problem. By creating a rich context and leveraging self-generated insights, the LLM can provide more detailed, accurate, and nuanced answers, showcasing the full extent of its capabilities. Table 11 presents the prompt templates utilized for Least-to-Most Prompting across all datasets in our experiments.
Dataset Prompt
BoolQ inference prompt: Based on the passage:"passage", answer True/False to the question: ’question’ using knowledge of the passage:"knowledge"
knowledge generation prompt: Generate Knowledge about the passage: "passage".
CSQA inference prompt: Choose the answer:"question", A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4",E. ‘‘option 5" using knowledge of the question:"knowledge"
knowledge generation prompt: Generate Knowledge about the question: "question".
IWSLT inference prompt: Translate "english text": to French using definitions of the keywords:"knowledge"
knowledge generatiuon prompt: Generate definitions in french of each word in the text: "english text".
SamSum inference prompt: Summarize the Dialogue: "dialogue" using the interpretation of the dialogue:"knowledge"
knowledge generation prompt: Generate interpretation about the dialogue: "dialogue".
Table 11: Prompt templates of different datasets for Generated Knowledge Prompting
B.6 Adaptive HPF
The prompt-selector in adaptive HPF selects the prompting level based on the task complexity to address the task. Llama 3 8B serves as the prompt-selector in the experiments. The prompt template was meticulously designed to ensure maximum clarity, aiming to reduce hallucinations and select the most effective prompting strategy.
Prompt Template: Choose the most effective prompting strategy among five available strategies for the task. Begin with the lowest indexed strategy and progress to higher indexed strategies if the earlier ones are not effective. For a given task, the prompting strategies are:
  • Role Prompting: Defines a role for the model in solving the task.
  • Zero-shot Chain of Thought prompting: Stimulates reasoning and problem-solving by including the phrase ’Let’s think step by step’ without offering previous examples related to the task.
  • Three-shot Chain of Thought prompting: Offers three examples related to the task to guide the model’s reasoning process.
  • Least-to-most prompting: Uses a sequential method to derive essential insights from the task in order to solve it.
  • Generated Knowledge Prompting: Integration and application of external knowledge to accomplish the task. The external knowledge is generated using some other model based on the task.
Select only the index (do not provide the name) of the most effective prompting strategy.
Figure 8: HP-Scores across all datasets and LLMs for manual HPF and Adaptive HPF.
Appendix C Human Annotation Policy
To ensure the quality of HP-Scores scored by human experts for the datasets, we implemented a rigorous human annotation process. Human annotators were tasked with calculating the HP-Score for each sample in a given dataset. The HP-Score relative to human experts is a measure of the expert capability in addressing the task based on criteria of HPT. Each sample was scored on a scale from 1 (Lowest prompting level) to 5 (Highest prompting level) for the following criteria:
C.1 Criteria
  1. 1.
    Basic Understanding and Reproduction
    • This criterion evaluates the annotator’s ability to comprehend and accurately reproduce the content. A high score indicates clear and precise reproduction of the information is required for solving the task.
  2. 2.
    Understanding and Interpretation
    • This criterion assesses the depth of understanding and the ability to interpret the information correctly. A high score indicates a detailed and precise understanding of the task is necessary to solve the task.
  3. 3.
    Analysis and Reasoning
    • This criterion measures the ability to analyze the information and apply logical reasoning. A high score demonstrates the need to analyze the information to solve the task.
  4. 4.
    Application of Knowledge and Execution
    • This criterion evaluates the practical application of knowledge and the execution of tasks based on the relevant knowledge. A high score indicates effective application and the requirement of knowledge to solve the task.
C.2 Scoring Process
Each sample of every dataset was scored individually by the annotators on a scale from 1 to 5, where:
  • 1 = Poor
  • 2 = Fair
  • 3 = Good
  • 4 = Very Good
  • 5 = Excellent
The final HP-Score for each dataset was calculated by taking the mean of the values from these four criteria. This approach ensures a comprehensive evaluation of each sample by considering multiple performance criteria.
Appendix D Additional Results
D.1 HP-Scores Interpretation
Figure 8 visualizes HP-Scores for different datasets across various LLMs, comparing manual and adaptive frameworks. Here’s a detailed summary:
For the BoolQ dataset, the manual HPF HP-Scores are generally lower across all LLMs compared to the Adaptive HPF. The adaptive scores are significantly higher, indicating better performance with the manual HPF for solving BoolQ.
For the CSQA dataset, manual HPF HP-Scores are moderate, slightly higher than those for BoolQ. However, similar to BoolQ, the Adaptive HPF HP-Scores are consistently higher, again showing the superiority of the manual HPF over adaptive HPF in solving CSQA.
In comparison to the other datasets, the manual HPF HP-Scores for the IWSLT (0.15) and IWSLT (0.20) datasets both show a significant increase. This increase is particularly noticeable for LLMs such as Phi 3 3.8B and Gemma 7B, which indicate a lack of ability to handle IWSLT. With comparable evaluation scores to manual HPF, which suggests no improvement in adaptive HPF, the adaptive HPF HP-Scores continue to be consistently high.
For both SamSum (0.15) and SamSum (0.20) datasets, the manual HPF HP-Scores are notably high for LLMs such as Phi 3 3.8B and Gemma 7B. Adaptive HPF HP-Scores are consistently high compared to manual HPF HP-Scores.
As evidenced by their lower HP-Scores, manual HPF routinely outperforms adaptive HPF in all tasks and LLMs, including BoolQ, CSQA, and SamSum, where it can outperform human experts. With the exception of solving SamSum (0.15) by Mistral 7B in Adaptive HPF, no LLM was able to complete any task with a HP-Score of less than 5, suggesting the need for additional prompting strategies in HPF or improved framework to avoid hallucinations in selecting prompting level by prompt-selector.
D.2 Hallucination in Adaptive HPF
Hallucinations in prompt-selector refer to instances where the LLM generates incorrect or misleading prompting levels, or nonsensical information that is not supported by HPF. These hallucinations can occur across various tasks, including question answering, multiple-choice questions, translation, and summarization. Below, we discuss the occurrence of hallucinations in the Llama 3 8B which was utilized as prompt-selector in our experiments across different tasks and iterations, highlighting how the LLM’s performance varies and where it tends to hallucinate.
As shown in Figure 9, for the BoolQ task, the prompt-selector initially struggles, indicated by the iterations where it reaches Level 4 with hallucinations (represented by ’…’). However, by the fourth iteration, Llama 3 8B manages to answer correctly at Level 2. For the CSQA task, prompt-selector exhibits hallucinations initially, shown by Level 4 and Level 0 (not included in HPF) responses. Eventually, it corrects itself by the third iteration, providing the correct answer at Level 2. For the IWSLT task, prompt-selector demonstrates a consistent pattern of hallucinations across multiple iterations. Even though Llama 3 8B attempts the translation at Level 2 multiple times, it ultimately fails to provide a correct translation, indicating a persistent hallucination. For the SamSum task, prompt-selector shows initial hallucinations in the first three iterations (Level 4). However, by the fourth and fifth iterations, the prompt-selector starts producing lower levels, finally, Llama 3 8B achieving correct answer at Level 2 in the last iteration .
The results indicate that Llama 3 8B exhibits hallucinations in selecting prompting levels across various tasks and iterations resulting in higher HP-Scores for Adaptive HPF, with performance varying significantly. While the LLM can eventually produce correct answers, as seen in the BoolQ and SamSum tasks, it often requires multiple attempts and may still fail in tasks like IWSLT translation.
Figure 9: Hallucination of prompt-selector in choosing the most effective prompting level to solve corresponding tasks by Llama 3 8B.
Appendix E Additional Experiments
E.1 Increasing Threshold for IWSLT and SamSum Datasets
When comparing the performance of various LLMs, Llama 3-8B stands out for its superior results on the IWSLT and SamSum tasks, consistently performing better in both manual and adaptive frameworks. At a threshold of 0.2, the performance of all LLMs reached saturation for these tasks, indicating that increasing the threshold further did not yield improvements. Both adaptive and manual frameworks struggled with thresholds higher than 0.2, failing to meet or exceed this benchmark for most samples within their iterations. Even the most capable LLM, Llama 3-8B, was unable to address the majority of samples across any prompting level within the HPF. This demonstrates the limitations of current HPFs at higher thresholds and highlights the ongoing challenges in achieving better performance in LLMs.
E.2 Adaptive HPF with Varying Prompt-selectors
Experiments demonstrated that Adaptive HPF is most effective when Llama 3 8B is used as a prompt-selector compared to LLMs, attributed to its ability to produce the lowest number of hallucinations when selecting the prompting level within the iterations of the adaptive HPF. Conversely, Gemma 7B, which exhibited numerous hallucinations across most tasks, was found to be the least effective adaptive HPF. Consequently, the choice of the prompt-selector emerges as a crucial factor influencing the efficiency of the adaptive HPF.